Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-190_2000-196

Mission Operations Days: 2000/190 to 2000/196
July 08, 2000 00:00:00 GMT to July 14, 2000 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
MODIS has an anomaly with the LWIR focal plane. See description below.
ComponentState(s)Comment
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On, #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOffCFPA Anomaly Investigation
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA On; B OffNominal
Flight SoftwareRev BD + 2 sets of patches
  • FR Band 14H Gain Change (10 commands)
  • FR Band 27 Gain Change (10 commands)
Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled66, 67Nominal

This Week’s Completed MODIS Activities:

Saturday, July 8, 2000

None

Sunday, July 9, 2000

None

Monday, July 10, 2000

2000/192 20:02:00 ATC - Blackbody to 270K

Tuesday, July 11, 2000

2000/193 14:02:00 ATC Blackbody to 315K
2000/193 14:00:00 Real-time Set Blackbody duty cycle to 100%
2000/193 21:00:00 ATC Blackbody to 270K
2000/193 21:30:00 Real-time Set Blackbody duty cycle to 33%

Wednesday, July 12, 2000

2000/194 14:40-14:45 Real-time Band 27 Gain changes from MCST
2000/194 14:45 Real-time Dump Gain Table
2000/194 19:00:00 ATC Blackbody to 290K
2000/194 21:00:00 – 21:37:13 ATC Load - OA25: SRCA 1W Spatial w/Sector Rotation, the BC and ND Filter
2000/194 22:45:00 – 23:22:13 ATC Load - OA25: SRCA 10W Spatial w/ Sector Rotation and Spectral Slit, the BC and ND Filter

Thursday, July 13, 2000

None

Friday, July 14, 2000

2000/196 16:15:50 Real-time Turn off LWIR Heater

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Saturday, July 15, 2000

None

Sunday, July 16, 2000

None

Monday, July 17, 2000

2000/199 ?? Real-time Open SVD 100 steps further
2000/199 20:02:00 ATC - Blackbody to 270K

Tuesday, July 18, 2000

200/14:00:00 ATC Blackbody to 315K
200/14:?? Real-time Set Blackbody duty cycle to 100%
200/21:00:00 ATC Blackbody to 270K
200/21:?? Real-time Set Blackbody duty cycle to 33%

Wednesday, July 19, 2000

201/19:00:00 ATC Blackbody to 290K

Thursday, July 20, 2000

None

Friday, July 21, 2000

Roll maneuver from 12:14:08 to 12:19:38 (not scheduled yet)

Maneuvers:

Friday, July 21, 2000 - MODIS Roll Maneuver (2000/203/12:07:22 – 12:25:24)

MODIS Anomalies:

The Radiative Cooler has negative margin, therefore the cold focal planes are no longer under thermal control. This affects bands 5, 6, 7, and 20-36. Analysis is underway. Solar Diffuser Door operations are suspended during this analysis.

Current maximum and minimum temperature ranges of the LWIR FPA:
Thursday, July 13, 2000:
2000/195 max = 83.859K
2000/195 min = 83.648K

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open.

The new gain values for Band 27 are as follows:
Detector 1 - 238
Detector 2 - 219
Detector 3 - 231
Detector 4 - 238
Detector 5 - 238
Detector 6 - 238
Detector 7 - 238
Detector 8 - 238
Detector 9 - 238
Detector 10 - 238

MODIS Telemetry Trends:

See MODIS Anomalies section.

Non-MODIS Significant Events:

Monday, July 17 (Day 199) the EOS Operations Center will be switching to the upgraded on-line operations software build, Rev C. They will also be upgrading to Project Database (PDB) 28. The MODIS IOT has reviewed these upgrades and is ready for the switchover.

Data processing by EDOS is currently running at a "degraded level". For a detailed summary of the events that have affected their products, view the EOS System Status web page (http://jupiter02.gsfc.nasa.gov:591/sysstat/) and select the EDOS summary link at the top of the page. A key item to note is from day 6/27/2000 at 12:11:19 PM: " The second major issue is the amount of data that EDOS is receiving since ASTER has been fully turned on." With the new, higher data output of ASTER, even less "spare" time will be available to EDOS for the processing of backlogged or reprocessing of data.

Limited Life Item Status:

SRCA 10W Lamp #1: 170.4 of 500 hours
SRCA 10W Lamp #2: 134.3 of 500 hours
SRCA 10W Lamp #3: 143.5 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 555.6 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1381 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements